on 1/5 a stock loses 25% of it's previous days value the next day 1/6 it gains 25% of the previous days value if the stock is worth 2500 on 1/6 what was it worth on 1/4?
Answer by xcentaur(357)

let the value of the stock be 'x'
Now on 1/5 it loses 25%
So now the stock is worth 75% of x = (75/100)*x = (3x/4)
Now on 1/6 the stock gains 25% of its value on 1/5
So now the stock is worth 125% of its value before
So now it is = 125% of (3x/4)
=(125/100)*(3x/4)
=(5/4)*(3/4)*x
This equals 2500 as given
Therefore,
2500 =(5/4)*(3/4)*x
x=(2500*4*4)/(5*3)
x=40000/15=2666.67


Hope this helps,
xC